OpenStudy (anonymous):

the square root od 18x^2

OpenStudy (anonymous):

\[\sqrt{18x ^{2}}\]

OpenStudy (anonymous):

\[3\sqrt{2}\]

OpenStudy (anonymous):

\[3\sqrt{2}x\]